Reacting with water in an acidic solution at a particular​ temperature, compound A decomposes into compounds B and C according to the law of uninhibited decay. An initial amount of 0.50 M of compound A decomposes to 0.47 M in 30 minutes. How much of compound A will remain after 2 ​hours? How long will it take until 0.10 M of compound A​ remains?
Answer by Boreal(15235)

A=Aoe^(kt)
0.47=0.50e(k*0.5), using hours as units
0.94=e^0.5k
ln both sides
-0.0619=0.5k
k=-0.1238
for 0.10 M
0.10=0.50e^(-0.1238*t)
0.2=e^(-0.1238t)
-1.61=-0.1238t
t=13.00 hours